OpenStudy (anonymous):

Which are counter examples of the following statement? If the sum of x and y is negative, then both x and y are negative. x = -10 and y = 5 x = -2 and y = -4 x = -6 and y = 0 x = 15 and y = -2

OpenStudy (michele_laino):

please what is: \[15+(-2)=...\]

OpenStudy (michele_laino):

hint: \[15+(-2)=15-2=...\]

Directrix (directrix):

You are looking for two numbers that add to a negative number BUT one of the numbers added is not negative.

Directrix (directrix):

Consider this option: x = 15 and y = -2 The sum of the two is number is 13 so they do not satisfy the hypothesis of having a negative sum. So, that option is out. @theviolentredhead

OpenStudy (anonymous):

B @Directrix

Directrix (directrix):

Not option B, @theviolentredhead Because -2 and -4 sum to -6 BUT they are both negative. We are looking for the sum of two numbers that is negative, but the numbers added to get that sum are not both negative.l

Directrix (directrix):

So, Options B and D have been ruled out, so far.

Directrix (directrix):

@theviolentredhead What is the sum of -10 and 5 ?

OpenStudy (mathmath333):

oh sry option \(b\) is in correct

OpenStudy (anonymous):

but is says i can have mroe then 1 answers

OpenStudy (mathmath333):

oh now it makes sense

OpenStudy (anonymous):

???

Directrix (directrix):

I know it. If you will just do as I ask, we will finish this problem in a minute and be gone. We cannot get give you the answer so please help us. Again, @theviolentredhead what is the sum of -10 and 5? (We are testing option A)

OpenStudy (anonymous):

so A and B ?

Directrix (directrix):

what is the sum of -10 and 5?

OpenStudy (mathmath333):

whay is \(\Huge -10+5\)

OpenStudy (mathmath333):

i mean *what is

OpenStudy (anonymous):

-5

Directrix (directrix):

Okay. So this option: x = -10 and y = 5 is correct because the two numbers add to a negative but one of them is not negative. Option A is correct and there is one more, you said.

Directrix (directrix):

Option C test: x = -6 and y = 0

Directrix (directrix):

What do you get when you add -6 and 0? What is -6 + 0 = ? @theviolentredhead

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

so A B and C are correct

OpenStudy (anonymous):

No, only A & C are correct

Directrix (directrix):

>>>so A B and C are correct NO

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

happy new year!!!!!

Directrix (directrix):

@dazz5 I do not appreciate your interrupting the thread. Now, tell us why C is correct.

OpenStudy (anonymous):

wait A and C are correct? and not b

OpenStudy (anonymous):

???

OpenStudy (mathmath333):

in option \(b\) the sum of -2 and -4 is negative but both-2 and -4 are negative so option b is not correct

OpenStudy (anonymous):

Sorry @Directrix I did not mean to interrupt, I was just trying to help. C is correct because -6 + 0 = -6 (so it evaluates to a negative answer and our individual x and y values are not both negative (i.e.one is positive, the other negative). Therefore it follows the rules of the counterargument .

OpenStudy (anonymous):

This goes for A, as stated above earlier, (-10 + 5= -5) The answer (-5) is negative and the actual x and y values have different signs.

Directrix (directrix):

@dazz5 No worries. On this, I wanted to comment: -6 + 0 = -6 (so it evaluates to a negative answer and our individual x and y values are not both negative (i.e.one is positive, the other negative). Zero is not a positive number. It is neither positive nor negative. Yet, option does qualify as a counterexample because the sum of -6 and 0 is negative and while -6 is negative, zero is not.